OpenStudy (anonymous):

Algebra 2 Help; g(f(x))

OpenStudy (anonymous):

My choices are... x x-3/8 8x+21 x+6

OpenStudy (anonymous):

@Compassionate Could you help with this, please? I've already asked this question but one of the suggestions wasn't an answer

OpenStudy (anonymous):

I believe it is "x".

OpenStudy (anonymous):

@Jhannybean I could use your help as well, please.

OpenStudy (jhannybean):

I don't know what your problem is though, to solve

OpenStudy (anonymous):

whoops, forgot something. f(x)=x-3/8 ; g(x)=8x+3 sorry

OpenStudy (compassionate):

g(f(x)) f(x)=x-3/8 ; g(x)=8x+3 First solve for f(x) then plug it into g(x)

OpenStudy (jhannybean):

\[\large g(f(x)) = 8(\frac{ x-3 }{ 8})+3\]

OpenStudy (jhannybean):

Now just simplify. \[g(f(x)) = x-3+3\] and you'll have your answer :)

OpenStudy (anonymous):

Oh, ok, so it is "x"? =D

OpenStudy (jhannybean):

mmhmm :)

OpenStudy (anonymous):

Awesome ^-^ Thank you, you two for coming.